OpenStudy (anonymous):

Problem:6 over 10%x45+____ Choices:A. .027B. .27C. 2.7D. 27

OpenStudy (anonymous):

this is the last one i will do for you. after two examples, you should be able to do it. refer back to either one if you need help. (6/10) % = .6 % = .1 % * 6 45 * .1 % = 45 * .1 * 1% = .1 * .45 = .045 .045 * 6 = .27

OpenStudy (anonymous):

math is my worst subject. You're truly the best ! thank's for helping me !